А es au aleu an ic d es s ring about 200 peop e on a F steaks on hand to meet customer demand? ay even and beneves hal abou 2 % o he patrons will o de he che s s ak speud How many o hoss meals should h plan on er 10de o be pre sure μ+2σ of having enough The restaurateur should have steaks on hand Round up to the nearest whole number)
Answer #1

Solution :

Given that,

p = 20% = 0.20

q = 1 - p = 1 - 0.20 = 0.80

n = 200

Using binomial distribution,

mean = n * p = 200 * 0.20 = 40

standard deviation = \sqrt n * p * q = \sqrt 200 * 0.20 * 0.80 = 5.6569

\mu + 2\sigma = 40 + 2 * 5.6569 = 51

The restauateur should have 51 steaks on hand.